OpenStudy (anonymous):

What is the equation of a line passing through the point (a, b) and having a slope of b?

OpenStudy (anonymous):

do u have a graph @sherlocked1270

OpenStudy (anonymous):

no that's the whole question

OpenStudy (anonymous):

y-b = b(x-a)

OpenStudy (anonymous):

\[y-b= b(x-a)\]

OpenStudy (anonymous):

they are correct

OpenStudy (anonymous):

can u explain

OpenStudy (anonymous):